# NIMO: a Nonlinear Interpretable MOdel

### v1

@@ B.3 Extension to sub-$\ell_{1}$ pseudo-norms
In some cases sparsity is a crucial aspect in the learning process. One of the limitations of lasso is the well-known over-shrinkage effect, which means that shrinkage of uninformative features could be achieved at the expense of shrinking the retained features as well. One way to avoid this is to use sub-$\ell_{1}$ pseudo-norms, which in the limist of the $\ell_{0}$ pseudo-norm achieve sparsity without any shrinkage. The main challenge with sub-$\ell_{1}$ pseudo-norm is that the resulting objective function becomes non-convex and harder to optimize. However, recent work has successfully used sub-$\ell_{1}$ norms as a prior on regression problems [Negri et al., 2023], which is similar to our setting.
@@
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (27)
In order to generalize to sub-$\ell_{1}$ pseudo-norms we need the following simple modification on $\nu_{i}$:
$\sum_{i=1}^{d}\frac{1}{\nu^{\delta}}=C,\quad 0<\delta\leq 1.$ (28)
We now showcase that we can follow a procedure completely similar to what we had with the original adaptive ridge regression. Fist, we reformulate the objective with the method of Lagrangian multipliers:
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}$ (29)
@@
Note that when $\delta=1$, we recover the Lasso $\ell_{1}$ norm. When $\delta\rightarrow 0$, it approaches the $\ell_{0}$ pseudo-norm, which basically counts the number of nonzero entries. We now apply the same re-parameterization:
The objective function then reads as
$\min_{\bm{\gamma}}\|\bm{y}-XD_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}c_{i}^{2\delta}=\lambda^{\delta}C,\quad c_{i}>0\;.$ (33)
For each fixed $\bm{c}$, $\bm{\gamma}$ can now be obtained in closed form:
$\hat{\bm{\gamma}}=\left(D_{\bm{c}}X^{T}XD_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}X^{T}\bm{y},$ (34)
and $\bm{\beta}$ can be computed as before by:
$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ (35)
Once we compute the $\hat{\bm{\beta}}$, we can substitute it back into the original objective. Taking the constraint into consideration leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ only:
$\min_{\bm{c}}\|\bm{y}-X\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}c_{i}^{2\delta},$ (36)
and we can optimize it through gradient descent.
@@ Appendix C NIMO: Implementation details
C.1 Model Formulation
We formulate NIMO as follow:
$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;.$ (37)
The architecture of our model involves two parts, as shown in Fig. 2. The left part is the nonlinear neural network, which operates on the input feature $\bm{x}$ element-wise. The right part is the linear model, with coefficients $\beta_{j}$ multiplied by the neural network output and then summed together.
In practice, each feature component $x_{j}$ of the input vector $\bm{x}$ is temporarily masked out to construct $\bm{x}_{-j}$, which is then passed through a neural network $g_{\bm{u}_{j}}$. The output of this network is added to a constant $1$ and then multiplied by the original feature value $x_{j}$, effectively producing a rescaled version of that feature:
$\tilde{x}_{j}=x_{j}\cdot(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))$ (38)
@@ C.2 Practical Implementation
We implement our model with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019]. While in principle each feature $x_{j}$ could have its own dedicated network $g_{\bm{u}_{j}}$, this can be computationally expensive for high-dimensional inputs. To reduce overhead, we instead use a shared neural network $g_{\bm{u}}$ across all components.
@@
To ensure the network remains aware of which feature is currently masked out, we append a positional encoding to $\bm{x}_{-j}$ before passing it through the shared network. We assign each masked input feature $\bm{x}_{-j}$ a unique binary vector corresponding to its index $j$, represented in fixed-length binary form using $\lfloor\log_{2}d\rfloor+1$ bits. This compact encoding allows the shared neural network to remain index-aware with minimal overhead.
@@
The shared network is implemented with 3 fully-connected (fc) layers. To encourage the network to select only certain relevant features contributing to the nonlinearity, we apply a group penalty on the first fc layer. However, due to the compensation effect from the subsequent layers, the sparsity is difficult to achieve by simply applying this penalty. To mitigate this issue, we inject noise into the output of the first layer by adding scaled Gaussian noise. Concretely, we draw noise from a normal distribution with mean 0 and variance 1, scale it by 0.2, and add it to the output. The noise-injected output is then passed through a tanh activation function before being fed into the subsequent layers.
@@
After the second fc layer, a sin activation function is applied to capture potentially repeating or high-frequency cyclic patterns in the data [Sitzmann et al., 2020]. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. We choose Adam [Kingma, 2014] as the default optimizer. For regression, we regularize the network also by restricting the output range within $[-1,1]$, which works well in practice. For classification, we use a less restrictive reduction to the range $[-1,3]$.
@@ Appendix D Experiments
We conduct experiments on both synthetic and real-world datasets. All experiments are conducted on a workstation equipped with a single user-grade NVIDIA GeForce RTX 2080 Ti GPU. The methods we compare with are linear model (Lasso for regression, Logistic Regression for classification), neural network, LassoNet [Lemhadri et al., 2021] and Contextual Lasso [Thompson et al., 2023] (for regression only). Here we provide the details about the experiment setups.
D.1 Implementation details / Hyperparameters
@@
We use scikit-learn [Pedregosa et al., 2011] to implement the Lasso and Logistic regression model. More specifically, we use LassoCV and LogisticRegressionCV with default 5-fold cross validation to select the best model. For LogisticRegressionCV, the $\ell_{1}$ penalty is applied.
@@
We implement the naive neural network (MLP) with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019]. For a fair comparison, the MLP has the same structure as the nonlinear part in our NIMO model, except no positional encoding. We use Adam [Kingma, 2014] as the default optimizer. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. Strong dropout regularization (p=0.6) is used to avoid overfitting.
@@
Contextual Lasso
We get the code of Contextual Lasso from the original Github repository  [^3] https://github.com/ryan-thompson/ContextualLasso.jl, which was implemented in Julia language. We follow the examples in the documentation and configure the model to have a structure and number of parameters comparable to our own model.
@@ D.2 Synthetic experiments: Setup
We create the synthetic datasets by first sampling the features, and then compute the targets according to different nonlinear patterns. Random Gaussian noises are added to the targets. The feature values are sampled uniformly between $-2$ and $2$. For logistic regression, we pass the targets through a logistic (sigmoid) function to get probabilities between 00 and $1$, and then sample the final binary labels using a Bernoulli distribution with those probabilities. In all the settings, we use $200$ training samples and $100$ samples each for validation and test.
@@
For all the settings, we generate $y$ first, and then generate labels through the following procedure:
@@
    $y=$ $+10\cdot x_{1}\cdot[1+2\cdot\tanh(2x_{2})+\sin(x_{4})]$ (45)
    $+20\cdot x_{2}\cdot[1+2\cdot\cos(2x_{1})]$
    $-20\cdot x_{3}\cdot[1+2\cdot\arctan(x_{2}x_{4})]$
    $+10\cdot x_{4}$
    $-10+\epsilon$
@@
    $y=$ $-20\cdot x_{1}\cdot[1+\tanh(x_{2}x_{4})]$ (46)
    $+20\cdot x_{2}\cdot\left[1+\frac{2}{\pi}\arctan(x_{4}-x_{5})\right]$
    $+30\cdot x_{4}\cdot\left[1+\tanh(x_{2}+\sin(x_{5}))\right]$
    $-10\cdot x_{5}\cdot\left[1+(2\sigma(x_{1}x_{4})-1)\right]$
@@ D.3 Synthetic experiments: Additional results
Regression
The advantage of our model is that it not only recovers the underlying linear coefficients but also selects sparse features for the neural network, thanks to the noise injection and group penalty. In the main paper, we presented the recovered linear coefficients. Here, we show the sparse features selected by the neural network in Figure 7. We observe that even though all features are input to the neural network, only a small subset of features is selected. Moreover, these selected features align with those involved in the nonlinearity computation in our settings, which demonstrates the model’s ability to effectively identify and focus on the most relevant features for the task.
@@ D.4 Real-dataset experiments
We conduct experiments on two real-world datasets: the diabetes dataset [Efron et al., 2004] and the Boston housing dataset [Belsley et al., 2005]. Each dataset is randomly split into training (60%), validation (20%), and test (20%) sets, repeated over 5 independent runs. For each split, we perform a grid search to select the optimal penalty parameter based on the validation loss (or validation accuracy for classification tasks). Error bars and confidence intervals are reported across these different splits.

### v2

@@ B.3 Extension to sub-$\ell_{1}$ pseudo-norms
In some cases sparsity is a crucial aspect in the learning process. One of the limitations of lasso is the well-known over-shrinkage effect, which means that shrinkage of uninformative features could be achieved at the expense of shrinking the retained features as well. One way to avoid this is to use sub-$\ell_{1}$ pseudo-norms, which in the limits of the $\ell_{0}$ pseudo-norm achieve sparsity without any shrinkage. The main challenge with sub-$\ell_{1}$ pseudo-norm is that the resulting objective function becomes non-convex and harder to optimize. However, recent work has successfully used sub-$\ell_{1}$ norms as a prior on regression problems (Negri et al., 2023), which is similar to our setting.
@@
$\min_{\bm{\beta}}\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (27)
In order to generalize to sub-$\ell_{1}$ pseudo-norms, we need the following simple modification on $\nu_{i}$:
$\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}=C,\quad 0<\delta\leq 1.$ (28)
We now apply the same re-parameterization:
The objective function then reads as
$\min_{\bm{\gamma},\bm{c}}\|\mathbf{y}-XD_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}c_{i}^{2\delta}=\lambda^{\delta}C,\quad c_{i}>0\;.$ (30)
Fixing $\bm{c}$, we can eliminate $\bm{\gamma}$ by expressing it in a closed-form in terms of $\bm{c}$:
$\hat{\bm{\gamma}}(\bm{c})=\left(D_{\bm{c}}X^{T}XD_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}X^{T}\mathbf{y}.$ (31)
Substituting it back into the original optimization problem and taking the constraint into consideration lead us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ only:
$\min_{\bm{c}}\|\mathbf{y}-XD_{\bm{c}}\hat{\bm{\gamma}}(\bm{c})\|^{2}+\lambda\|\hat{\bm{\gamma}}(\bm{c})\|+\mu\sum_{i=1}^{d}c_{i}^{2\delta},$ (32)
and we can optimize it through gradient descent.
To verify that this constraint leads sub-$\ell_{1}$ pseudo-norm to the original $\bm{\beta}$, first we reformulate the original objective with the method of Lagrangian multipliers:
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}$ (33)
@@
Note that when $\delta=1$, we recover the Lasso $\ell_{1}$ norm. When $\delta\rightarrow 0$, it approaches the $\ell_{0}$ pseudo-norm, which basically counts the number of nonzero entries.
@@ Appendix C NIMO: Implementation details
In this section, we provide a detailed introduction to the structure of NIMO, along with the strategies we used to improve training.
C.1 Model Definition
We define NIMO as follow:
$f(\mathbf{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j}))\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;.$ (36)
The architecture of our model consists of two components, as illustrated in Fig. 2. The left component is a nonlinear neural network that operates on the input features $\mathbf{x}_{-j}$, where the $j$-th element has been masked out. The right component is a linear model, in which coefficients $\beta_{j}$ are multiplied by the corresponding neural network outputs and subsequently summed.
In practice, for each feature component $x_{j}$ of the input vector $\mathbf{x}$, the value is temporarily masked to form $\mathbf{x}_{-j}$, which is then passed through a neural network $g_{\bm{u}_{j}}$. The output of this network is added to a constant $1$ and multiplied by the original feature value $x_{j}$, resulting in a rescaled version of the feature:
$\tilde{x}_{j}=x_{j}\cdot(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j}))$ (37)
@@ C.2 Practical Implementation
We implement our model with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). While the model definition is simple, in practice, several strategies are needed to make the training more efficient and effective.
@@
In principle each feature $x_{j}$ could have its own dedicated network $g_{\bm{u}_{j}}$, but this can be computationally expensive for high-dimensional inputs. To reduce overhead, we instead use a shared neural network $g_{\bm{u}}$ across all components. To ensure the network remains aware of which feature is currently masked out, we append a positional encoding to $\mathbf{x}_{-j}$ before passing it through the shared network. We assign each masked input feature $\mathbf{x}_{-j}$ a unique binary vector corresponding to its index $j$, represented in fixed-length binary form using $\lfloor\log_{2}d\rfloor+1$ bits. This compact encoding allows the shared neural network to remain index-aware with minimal overhead.
@@
In the main paper we showed that linear coefficients $\bm{\beta}$ in NIMO are interpretable as in a linear model. Even though the neural network contribution is not interpretable per se, we can still identify relevant nonlinear interactions by means of sparsity. The idea is to force the first layer of the neural network to be sparse, hence to use only relevant features. To do so, we apply a group penalty on the weight matrix $W\in\mathbb{R}^{p\times d}$ of the first fully-connected (fc) layer, where $d$ is the number of input features and $p$ is the hidden dimension of the first fc layer. In order to obtain sparsity at the input feature level, we need to consider neurons connected to the same input feature and penalize them as a group:
$\ell_{\text{group}}=\lambda_{\text{group}}\sum_{j=1}^{d}\|\mathbf{w}_{j}\|_{2}\;,$ (38)
where $\mathbf{w}_{j}=W[:,j]\in\mathbb{R}^{p}$ is the weight vector acting on the $j$-th feature. With this penalty on the first fc layer, we encourage it to select only certain input features. However, since an MLP computes a composition of functions, the subsequent fc layers can still compensate for the sparsity induced by the regularization in the first layer. Specifically, even if the first layer produces sparse outputs, subsequent layers can reweigh and transform these outputs to restore their influence on the final prediction. To mitigate this compensation effect, we inject noise into the output of the first layer. By doing so, we disrupt the deterministic signal path, making it harder for the subsequent layers to simply exploit the resulting weak, regularized, and noisy signal. As a result, the network is encouraged to focus on strong and robust signals associated with informative input features. This approach works very well in practice and allows to easily achieve sparsity in the neural network.
@@
After the second fc layer, a sin activation function is applied to capture potentially repeating or high-frequency cyclic patterns in the data (Sitzmann et al., 2020). The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. We choose Adam (Kingma, 2014) as the default optimizer. In all experiments, the lasso and group penalties are selected by performing a grid search over a predefined set of candidate values.
@@ Appendix D Experiments
We conduct experiments on both synthetic and real-world datasets. All experiments are conducted on a workstation equipped with a single user-grade NVIDIA GeForce RTX 5080 GPU. The methods we compare with are linear model (Lasso for regression, Logistic Regression for classification), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024).
D.1 Methods implementation
@@
We use scikit-learn (Pedregosa et al., 2011) to implement the Lasso and Logistic regression model. More specifically, we use LassoCV and LogisticRegressionCV with default 5-fold cross validation to select the best model. For LogisticRegressionCV, the $\ell_{1}$ penalty is applied.
@@
We implement the naive neural network (NN) with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). For a fair comparison, the NN has the same structure as the nonlinear part in our NIMO model, except no positional encoding. We use Adam (Kingma, 2014) as the default optimizer. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. Strong dropout regularization (p=0.6) is used to avoid overfitting.
@@
NAM
The official NAM implementation was in Tensorflow. We chose a well-maintained PyTorch implementation from the GitHub repository  [^3] https://github.com/lemeln/nam. This implementation provides a scikit-learn style interface, which can be easily used.
IMN
We use the original PyTorch implementation of IMN from their GitHub repository  [^4] https://github.com/ArlindKadra/IMN. All the arguments are kept at their default values except modifying the data loading logic to fit our datasets.
For all the methods above, we tried our best to make sure that they achieve reasonably good and consistent performance on all the datasets.
@@ D.2 Synthetic experiments: Setup
We create the synthetic datasets by first sampling the features, and then compute the targets according to different nonlinear patterns. Random Gaussian noises are added to the targets. The feature values are sampled uniformly between 2-2 and $2$. For logistic regression, we pass the targets through a logistic (sigmoid) function to get probabilities between 0 and $1$, and then sample the final binary labels using a Bernoulli distribution with those probabilities. In all the settings, we use $200$ training samples and $100$ samples each for validation and test.
@@
In order to create data for synthetic classification we follow a similar procedure as in the regression cases, where we directly control the linear coefficients and the nonlinear interactions. The only difference is that we further apply a link function and binarize the output and get the labels. Note that this way we loose information about the original linear coefficients, which cannot be recovered exactly anymore. The sparsity in the coefficients still remains crucial. For all the settings, we generate $y$ first, and then generate labels through the following procedure:
@@
[…]
@@
[…]
@@ D.3 Synthetic experiments: Additional results
Regression
As we mentioned in the main paper, we further apply group $\ell_{2}$ regularization to the weight matrix of the first fully connected (fc) layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions. We provide the details in Appendix C.2. Here, we show the sparse features selected by the neural network in Figure 8. We observe that even though all features are input to the neural network, only a small subset of features are selected. Moreover, these selected features align with those involved in the nonlinearity computation in our settings, which demonstrates the model’s ability to effectively identify and focus on the most relevant features for the task.
@@
Our method can be extended to generalized linear models and here we showcase it for logistic regression. Also in this case we show various settings for different dimensionality, nonlinear terms and different number of uninformative features. In Table 3 we report the classification accuracy on the test set for all methods. Overall, our model significantly outperforms logistic regression on all settings. When analyzing the learned coefficient, we can see that only NIMO recovers the correct sparsity patterns. As argued before, this allows to learn the correct nonlinear interactions. Other nonlinear models and hybrid approaches perform similarly to NIMO but fail to provide the interpretability of the coefficients as in a linear model.
Table 3: Classification accuracy for synthetic classification settings. We compare with logistic regression, NN and LassoNet. In each setting we use 200 samples.
Features Log. Regr. NN NIMO LassoNet
Setting 1 3 0.59 0.90 0.92 0.89
Setting 2 10 0.68 0.83 0.85 0.82
Setting 3 50 0.74 0.69 0.84 0.90
@@ D.4 Real-dataset experiments
We conduct experiments on three real-world datasets: the diabetes dataset (Efron et al., 2004), the Boston housing dataset (Belsley et al., 2005) and the superconductivity dataset (Hamidieh, 2018). Each dataset is randomly split into training (60%), validation (20%), and test (20%) sets, repeated over 5 independent runs. For each split, we perform a grid search to select the optimal penalty parameter based on the validation loss (or validation accuracy for classification tasks). Error bars and confidence intervals are reported across these different splits. The main results are shown in Figure 5.
Diabetes dataset
Figure 11: Diabetes datasets results. Left: MSE loss on test set. All methods achieve comparable results. Middle: Learned linear coefficients by Lasso and NIMO. NIMO selects sparser coefficients (“s1, “s2”, “s4”). Right: the first fc layer of NIMO is very sparse; few nonlinearities are relevant.
Boston dataset
The Boston housing dataset contains 506 instances, each with 13 features, and the target is the median value of owner-occupied homes. As shown in Figure 12 (left), our method performs significantly better than Lasso regression and shows compatible MSE test loss with the other nonlinear and hybrid approaches. This suggests that nonlinear interactions are crucial to explain the data. While the other methods are able to capture those nonlinearities, they lack the interpretability provided by our model. It is important to note that the Boston housing dataset has been criticized for containing controversial features  [^5] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html. The feature of interest is termed “B”. From the coefficient plot in Figure 12 (middle), we observe that the significance of the feature “B” is greatly reduced in our model compared to Lasso regression. This surprising result suggests that the role of the feature “B” towards the prediction might be irrelevant. Note that the feature “B” is also not relevant for the nonlinear terms, as shown in Figure 12 (right).
Figure 12: Boston datasets results. Left: Comparison of MSE losses. NIMO performs on par with other methods and significantly outperforms Lasso, suggesting nonlinear interactions are crucial. Middle: Learned linear coefficients by Lasso and NIMO. The two models learn coefficient with a comparable sparsity. Relevantly, the feature “B” is almost irrelevant for NIMO. Right: sparsity in the first fc layer of NIMO.
Superconductivity
The superconductivity dataset contains 21,263 instances, each with 81 features. The target is to predict the critical temperature based on the features extracted.
Figure 13: Superconductivity dataset results. Left: Comparison of MSE losses. Right: sparsity in the first fc layer of NIMO.
In Figure 13, we observe that NIMO achieves the best performance. With respect to sparsity in the first fc layer, the presence of strong correlations among several of the 81 features makes it challenging to obtain clear sparsity patterns. Nevertheless, certain patterns can still be discerned (Figure 13, right). Figure 14 shows the coefficients for Lasso and NIMO on this dataset.
Figure 14: Coefficients for Lasso and NIMO on superconductivity dataset.
D.5 Ablation study of feature masking
To verify the effectiveness of masking out the $j$-th component of an instance when passing it through the neural network, we conducted an ablation study. Specifically, we examined the effect of masking out the $j$-th component as well as the effect of removing position encoding. From the ablation results across the three settings (Figure 7, Figure 15, and Figure 16), we observe several consistent patterns:
  • Without masking the $j$-th component, the first fc layer of the neural network can still exhibit sparsity patterns, but they are less distinct than when masking is applied. We attribute this to the position encoding, which enables the network to identify the feature index.
  • Without masking the $j$-th component, the magnitudes of the linear coefficients are distorted, reducing the interpretability of the marginal effects at the mean (MEM).
  • Further removing the position encoding distorts the sparsity patterns in the first fc layer and substantially alters the magnitude of the linear coefficients, thereby undermining interpretability.
Figure 15: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 2.
Figure 16: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 3. Coefficients are displayed for the first 10 features.
Appendix E LLM statement
ChatGPT is used only for polishing the sentences.